ILK has preferred stock selling for 97 percent of par that pays an 8 percent annual coupon. What would be ILK’s component cost of preferred stock? (Round your answer to 2 decimal places.)
Answer #1

Information provided:

Par value= $100

Current price= 0.97*1,00= $97

Annual coupon= 8%

Coupon payment= 0.08*100= $8

Cost of preferred stock is calculated using the below formula:

= Preferred dividend/ Market price of preferred stock

= $8/ $97

= 0.0825*100

= 8.25%.

In case of any query, kindly comment on the solution.